OpenStudy (anonymous):

When you solve the equation x plus nine equals fifteen, how do you know what number to subtract from both sides? Why is it important to subtract the same number from both sides of the equation?

OpenStudy (anonymous):

x+9=15 When you look at the equation above, there is one variable you need to solve for, x. To solve for it, you need to subtract 9 from 15 in order for it to be all alone on the left side. visually, it looks like this. The parenthesis indicate the original equation you were given. (x+9)-9=(15)-9 therefore, x+0=15-9 x=6 Its important to subtract the same number from both sides because if you dont, you wont get the right answer. Sorry if that isn't very helpful. Just remember, whatever you do to one side, you must do to the other.